In a certain city a birth occurs on the average every 24 minutes and a death every half hour. A resident moves out of the city every 1.5 hours and a new person moves into the city every 4.5 hours. How long does it take on the average for the population to increase by 1 person? Please help me!
Answer by stanbon(75887)

In a certain city a birth occurs on the average every 24 minutes and a death every half hour. A resident moves out of the city every 1.5 hours and a new person moves into the city every 4.5 hours. How long does it take on the average for the population to increase by 1 person? Please help me!
-----
How many births in a day? 60
How many arrive in a day? 5 1/3
-----
Total in 24 hrs: 60 1/3
====================================
How many deaths in a day? 48
How many leave in a day? 16
-----
Total in 24 hrs: 64
-----
So you have a gain of 1 1/3 in 24 hrs.
----
When do you have a gain of 1 person?
Solve (4/3)/24 = 1/x
(4/3)x = 24
x = 18 hrs.
---
Every 18 hrs. the town gains one person.
=============================================
Cheers,
Stan H.